# PicoAudio2: Temporal Controllable Text-to-Audio Generation with Natural Language Description

### v1

@@ Abstract
Controllable text-to-audio generation (TTA) has attracted much attention recently. Although existing works can achieve fine-grained controllability based on timestamp information, sound event categories are limited to a fixed set. Moreover, since only simulated data is used for training, the generated audio quality and generalization performance on real data are limited. To tackle this issue, we propose PicoAudio2, improving temporal-controllable TTA via a new data processing pipeline and model architecture. Specifically, we use a grounding model to annotate event timestamps of real audio-text datasets to curate temporally-strong real data, in addition to simulation data from existing works. The model is trained on the combination of real and simulation data. Moreover, following PicoAudio, we encode timestamp information into a timestamp matrix to provide extra fine-grained time-aligned information to the model, on top of the coarse-grained textual description. Experiments show that PicoAudio2 exhibits superior performance in terms of temporal controllability and audio quality. The demo page can be accessed at https://HiRookie9.github.io/PicoAudio2-Page.
@@ 1 Introduction
Recently, text-to-audio generation (TTA) has received a lot of attention. TTA has broad application potentials in fields such as virtual reality and social media content creation. With the emergence of advanced generative models [1, 2], TTA has made remarkable progress [3, 4, 5, 6, 7]. Mainstream TTA models are able to generate high-fidelity audio from the given text description. The control of additional condition is also explored [8].
In TTA, the temporal controllability is crucial for enabling precise alignment between text description and sound events. Due to the lack of fine-grained temporal alignment between text and audio in common datasets [9, 10, 11], as well as the limited ability of text encoders to capture temporal information, current mainstream TTA models struggle to generate audio recordings that are precisely aligned with the input text.
Few works have investigated fine-grained controllability in TTA. Make-An-Audio 2 (MAA2) [6] leverages large language models (LLMs) to convert input text into structured descriptions specifying temporal occurrence (e.g., an event happens throughout, at the beginning, or at the end). However, such control remains coarse-grained. AudioComposer [12] also employs structured text to encode control information such as temporal occurrence and energy, achieving high control accuracy. However, captions in the training data are all simulated, resulting in limited audio quality when evaluated on real datasets(human-labelled real-world datasets, such as AudioCaps [9]), which will be shown in Section 5.1. PicoAudio introduces a timestamp matrix to indicate the occurrence time of events [13], achieving high audio quality and temporal controllability of pre-defined short-duration events. However, as the timestamp matrix is limited to pre-defined categories, it lacks the flexibility to accommodate natural language descriptions and open-ended event specifications.
@@
Fig. 1: The data curation pipeline. The left part shows the real dataset processing pipeline, where the TAG model extracts event timestamps and filter out data without omissions or overlaps. The right part shows the data simulation pipeline, where multi-event audio is simulated from preprocessed single-event segments with precise timestamp information. Captions are obtained by concatenating single-event descriptions.
To achieve precise temporal controllability and high audio quality in TTA, we propose PicoAudio2, as shown in table 1. Compared with PicoAudio, PicoAudio2 incorporates improvements in both data and model design, supporting free-text input. In terms of data, we propose data processing pipelines for both simulation and real data. For simulation data, we synthesize audio-caption-timestamp triplets using the method proposed in AudioTime [14]. Since the captions for individual events are category labels (not desired natural text), we convert these labels into textual descriptions. via LLMs and the Contrastive Language-Audio Pre-training (CLAP) model [15]. However, there is still a distribution gap between simulated datasets and real datasets, both in the audio and caption domains. Therefore, we incorporate real datasets into training to mitigate the gap. For real datasets, we segment captions into single-event descriptions using LLMs, and extract timestamp information with a Text-to-Audio Grounding (TAG) model [16].
In terms of model design, PicoAudio2 employs free-text descriptions for individual events to construct a timestamp matrix, thereby overcoming the limitations of fixed categories in PicoAudio. Leveraging this timestamp matrix, PicoAudio2 enables precise temporal control from natural language input. The matrix offers fine-grained timing information, while the original coarse caption supplies overall audio semantics. It also supports training with data that lacks timestamps.
@@
  • We propose a novel data processing pipeline, which enhances the simulated dataset and incorporates real data into training.
@@ 2 Temporally-Aligned Data Curation
To leverage a larger amount of data for training, we design curation pipelines for both simulation and real data, as shown in Figure 1. Typical real datasets, like AudioCaps [9], are structured as audio-caption pairs. Since these captions lack exact timestamps of sound events, we denote them as temporal coarse captions (TCC), and define such audio–caption pairs as temporally weak data. The pipelines are designed to generate temporally detailed captions (TDC), which contain timestamps for each sound event, and to construct audio–TCC–TDC triplets, thereby yielding temporally strong data for training.
@@ 2.1 Simulation Data
We adopt the approach in AudioTime [14] to synthesize high-quality audio-TCC-timestamp triplets. However, the single-event descriptions in TCC are category labels (e.g., ‘barking’, ‘speaking’). Given the gap between category labels and free-form text, we leverage large language models (LLMs) [17] to generate multiple free-text descriptions for each category. Additionally, we use LLMs to extract single-event descriptions from AudioCaps caption corpus. Then we assign category labels to these descriptions based on CLAP [15] similarities. For each category, we first select the top 30 free-text candidates based on CLAP scores, followed by manual review to ensure quality.
Ultimately, each category label is mapped to 15–30 text descriptions. For instance, “dog” is mapped to descriptions such as “a dog is barking and yipping” and “barking from a dog”. During training, category labels in TCC are mapped to randomly sampled descriptions from their corresponding candidate lists. These descriptions are then combined with timestamps to form TDC. Finally, we construct the audio-TCC-TDC triplets, which serve as temporally strong data.
@@ 2.2 Real Data
For real data where audio and TCC are already available, we first use LLMs to extract single-event descriptions from TCC. Then, the TAG model [16] is employed to detect the timestamps of corresponding events, thereby associating single-event descriptions with timestamps to form TDC. However, we observe that the TAG model is still limited in some scenarios: 1) inaccuracy arises when multiple events overlap; 2) the model may fail to detect any occurrences of event descriptions in TCC in certain audio clips. Therefore, to ensure timestamp accuracy, we filter out samples with timestamp overlaps or omissions, retaining only well-curated data as temporally strong data. In addition, the original datasets are also treated as temporally coarse data, with only the audio and corresponding TCC (i.e., original captions) used during training.
@@ 3 PicoAudio2
The overall structure of PicoAudio2 is shown in Figure 2. PicoAudio2 comprises a Variational Autoencoder (VAE), a text encoder, and a timestamp encoder. The VAE operates on raw waveforms to compress audio signals into latents $\mathbf{A}$ during training and transforms latents back to audio waveforms during inference. We employ the VAE in EzAudio [7]. The timestamp encoder extracts temporal information from TDC, obtaining timestamp matrix $\mathbf{T}$. The Flan-T5 [18] text encoder extracts semantic information $\mathbf{C}$ from TCC. VAE and Flan-T5 are frozen during training.
Fig. 2: PicoAudio2 framework. The red arrow represents the training process while the blue represents inference. During inference, the users can either provide detailed timestamps for each events, or a coarse description for LLM to infer the timestamp information.
@@ 3.1 Timestamp Encoder
The timestamp encoder extracts the timestamp matrix $\mathbf{T}$ from TDC to provide temporal information. For temporally coarse data (which lack TDC), $\mathbf{T}$ is set as a fixed embedding sequence. This design enables PicoAudio2 to generate valid audio with only TCC input. For temporally detailed data, the timestamp encoder first transforms single event descriptions in TDC to event-level features $\mathbf{a}\in\mathbb{R}^{C}$ using Flan-T5. Then these features are aggregated according to timestamps of each event, resulting in $\mathbf{T}\in\mathbb{R}^{T\times C}$:
@@
For each timestep $t$, features of occurring events are summed as $\mathbf{T}_{t}$. Compared with PicoAudio [13], which maps text descriptions to fixed classes, PicoAudio2 is capable of handling any events described by natural language while still providing the representative timestamp matrix $\mathbf{T}$ that remains temporally aligned with the audio latent $\mathbf{A}$.
During training, TDC are obtained by simulation or pre-processing real data. During inference, users can provide TDC directly, or provide TCC and rely on LLMs to transform TDC to TCC. Since LLMs are sufficiently powerful to generate reasonable TDC from TCC, we focus on guiding the model to generate audio recordings that adhere to provided TDC.
@@ 4 Experiments
Table 2: General audio quality and temporal controllability results. Best results are in bold.
@@
PicoAudio2 38.234 2.614 11.348 0.361 3.29 0.814 0.673 4.15
AudioCaps AudioLDM2 [3] 22.598 2.252 9.914 0.334 2.93
Tango2 [5] 36.301 2.304 9.886 0.392 3.38
MAA2 [6] 26.699 2.335 10.392 0.372 3.23
AudioComposer [12] 27.418 2.516 7.864 0.285 2.36
PicoAudio2 (w/o T) 30.347 2.460 12.373 0.352 2.75
PicoAudio2 31.540 2.547 12.256 0.373 3.18
4.1 Datasets
Both simulation and real data are used for training. For simulation data, we generate 64K audio clips. Each clip lasts at most 10 seconds and includes one to four events. For real data, we transform captions in AudioCaps [9] and WavCaps-ASSL subset [11] to TDC, resulting in 49K high-quality audio-TDC pairs. We also utilize around 106K temporally weak data samples from these two datasets. Overall, the total training data consists of two parts: ~106K temporally weak data (in the form of audio-TCC pairs) and ~113K temporally strong data (in the form of audio-TCC-TDC triplets). During training a sampling ratio of 1:2 between temporally weak and strong data is used.
4.2 Experimental Setup
The time resolution of the timestamp matrix is set to 20 ms. The diffusion backbone follows EzAudio [7], with 24 layers, 16 attention heads and a hidden size of 1024. PicoAudio2 is trained for 40 epochs with a maximum learning rate of $1\times 10^{-4}$. The learning rate takes a linear decay schedule with a weight decay of $1\times 10^{-6}$. During inference, the classifier-free guidance scale is set to 7.5.
4.3 Evaluation
The evaluation is done in terms of general audio quality and temporal controllability. For general audio quality, Frechet Distance (FD), Inception Score (IS), Kullback–Leibler divergence (KL) and CLAP score are used. For temporal controllability, we annotate the timestamps of each event using TAG [16]. As described in Section 2.2, we retain data without timestamp omission or overlap, naming this subset AudioCaps-DisJoint (DJ). We use Segment-F11 (Seg-F1[21] as the metric. Since the precise temporal control of multiple events is more difficult than single events, we filter AudioCaps-DJ to construct a multi-event subset (exclusively for temporal controllability evaluation), reporting the corresponding Seg-F11-ME.
Besides objective metrics, we report the Mean Opinion Score (MOS): MOS-Q for audio quality and MOS-C for temporal controllability. 10 high-educated raters without hearing loss are invited to score 10 random samples from AudioCaps and AudioCaps-DJ.
@@ 5.1 Generation Performance
We first compare PicoAudio2 with baselines, including AudioLDM2 [3], Tango2 [5], MAA2 [6] and AudioComposer [12], on general audio quality. Among them, MAA2 and AudioComposer require specified input formats to achieve temporal control so the inputs for these two models are constructed accordingly. For Tango2 and AudioLDM2, we input TCC and TDC respectively, and take the better-performing result for evaluation. Results are shown in Table 2. Per subjective and objective metrics, PicoAudio2 achieves performance comparable to leading models on both AudioCaps and AudioCaps-DJ, indicating high audio quality. On AudioCaps-DJ, AudioComposer performs notably worse. It takes both event and temporal information within its input descriptions, which may lead to suboptimal event fidelity. On AudioCaps, AudioComposer achieves better audio quality without timestamp input, further validating the negative influence of temporal information. By decoupling captions with temporal information, PicoAudio2 is capable of generating high-quality audio with either TCC or TDC inputs.
@@ 5.2 Ablation Studies
To verify the necessity of model design and data proposed in PicoAudio2, we further explore the effect of timestamp matrix and real training data. The timestamp matrix is first excluded from PicoAudio2, denoted as ‘w/o T’. This ensembles AudioComposer in that temporal information is embedded in the caption so the model relies solely on cross attention with the caption feature to encode temporal information. Results in Table 2 demonstrate that the exclusion of the timestamp matrix results in inferior temporal controllability, while still maintaining reasonable audio quality, validating the effectiveness of the timestamp matrix.
@@ 6 Conclusion
Mainstream TTA models struggle to achieve high audio quality and precise temporal control on natural language description. To this end, we propose PicoAudio2 in this work, incorporating new data processing pipelines and framework designs to improve temporal controllable TTA. In terms of data, PicoAudio2 designs pipelines for both simulation and real data, obtaining audio-TCC-TDC triplets for training. In terms of model, PicoAudio2 adopts a new timestamp matrix to represent event captions and their temporal occurrence to provide time-aligned feature for controllable generation. Objective and subjective evaluation show that PicoAudio2 achieves superior temporal controllability with audio quality comparable to mainstream TTA models. Due to challenges in real data annotation, PicoAudio2 train only on disjoint subsets of real datasets. It results in limited temporal control on overlapping events, which serves as a direction for our future work.

### v2

@@ Abstract
While recent work in controllable text-to-audio (TTA) generation has achieved fine-grained control through timestamp conditioning, its scope remains limited by audio quality and input format. These models often suffer from poor audio quality in real datasets due to sole reliance on synthetic data. Moreover, some models are constrained to a closed vocabulary of sound events, preventing them from controlling audio generation for open-ended, free-text queries. This paper introduces PicoAudio2, a framework that advances temporal-controllable TTA by mitigating these data and architectural limitations. Specifically, we use a grounding model to annotate event timestamps of real audio-text datasets to curate temporally-strong real data, in addition to simulation data from existing works. The model is trained on the combination of real and simulation data. Moreover, we propose an enhanced architecture that integrates the fine-grained information from a timestamp matrix with coarse-grained free-text input. Experiments show that PicoAudio2 exhibits superior performance in terms of temporal controllability and audio quality. The demo page can be accessed at https://HiRookie9.github.io/PicoAudio2-Page.
@@ 1 Introduction
Recently, text-to-audio generation (TTA) has garnered significant research interest due to its broad application in fields such as virtual reality and social media content creation. With the emergence of advanced generative models [1, 2], TTA has made progress in generating high-fidelity audio from textual descriptions [3, 4, 5, 6, 7]. Despite these improvements, fine-grained temporal control remains a critical and understudied challenge. Unlike text-to-speech tasks, where alignment is naturally structured, general audio datasets lack fine-grained temporal annotations [8, 9, 10], and text encoders often fail to capture precise temporal cues. As a result, mainstream TTA models struggle to generate audio recordings that are precisely aligned with the input text. Several recent works have begun exploring temporal controllability in TTA. For example, Make-An-Audio 2 (MAA2) [6] leverages large language models (LLMs) to convert free-text input into structured temporal descriptions (e.g., an event happens throughout, at the beginning, or at the end). However, such control remains coarse-grained. AudioComposer [11] also employs structured text to encode control information such as temporal occurrence and energy and achieves high control accuracy. However, captions in the training data are all simulated, resulting in limited audio quality on real-world datasets like AudioCaps [8], which will be shown in Section 5.1. PicoAudio [12] introduces a timestamp matrix to indicate the occurrence time of events, achieving high audio quality and temporal controllability of pre-defined short-duration events. However, as the timestamp matrix is limited to pre-defined categories, it lacks the flexibility to accommodate natural language descriptions and open-ended event specifications.
@@
Fig. 1: The data curation pipeline. The left part shows the real dataset processing pipeline, where the TAG model extracts event timestamps and data with omissions or overlaps are excluded. The right part shows the data simulation pipeline, where multi-event audio is simulated from preprocessed single-event segments with precise timestamp information. Captions are obtained by concatenating single-event descriptions.
To achieve precise temporal controllability and high audio quality given free-text input, we propose PicoAudio2, as shown in Table 1. PicoAudio2 incorporates improvements in both data and model design, supporting temporal control with free-text input. In terms of data, we develop dedicated data processing pipelines for both simulation and real data. For simulation data, we synthesize audio-caption-timestamp triplets following the AudioTime [13] method. Since the original event labels are categorical rather than natural language descriptions, we convert these labels into textual descriptions via LLMs and extract embeddings with the Contrastive Language-Audio Pre-training (CLAP) model [14]. However, there is still a distribution gap between simulated datasets and real datasets, both in the audio and caption domains. Therefore, we incorporate real datasets into training to mitigate the gap. For real datasets, we segment captions into single-event descriptions using LLMs, and extract timestamp information with a Text-to-Audio Grounding (TAG) model [15].
In terms of model design, PicoAudio2 employs free-text descriptions for individual events to construct a timestamp matrix. Single-event descriptions are first transformed to embeddings, which are then combined with corresponding timestamps to form the matrix. This matrix provides fine-grained temporal control signals, while the original coarse-grained caption offers global semantics. The integration of both inputs allows the model to achieve precise temporal alignment from open-ended language descriptions. In addition, such a design retains training flexibility and allows for the inclusion of data without timestamp annotations by replacing the matrix with a fixed embedding.
@@
  • We propose a novel data processing pipeline that enhances simulated dataset and incorporates real data into training.
@@ 2 Temporally-Aligned Data Curation
To leverage a larger amount of data for training, we design curation pipelines for both simulation and real data, as shown in Figure 1. Typical real datasets such as AudioCaps [8], consist of audio-caption pairs without precise temporal annotations. We refer to such captions as temporal coarse captions (TCC), and the corresponding audio–caption pairs as temporally weak data. Our pipelines are designed to convert them into temporally detailed captions (TDC), containing exact timestamps for each sound event, resulting in audio–TCC–TDC triplets that constitute temporally strong data for training.
@@ 2.1 Simulation Data
We adopt the approach in AudioTime [13] to synthesize high-quality audio-TCC-timestamp triplets. However, the single-event descriptions in TCC are categorical labels (e.g., ‘barking’, ‘speaking’). Given the gap between category labels and free-form text, we leverage LLMs [16] to generate multiple free-text descriptions for each category. Additionally, we use LLMs to extract single-event descriptions from AudioCaps caption corpus. Then we assign category labels to these descriptions based on CLAP [14] similarities. For each category, we first select the top 30 free-text candidates based on CLAP scores, followed by manual review to ensure quality.
Ultimately, each category label is mapped to 15–30 textual descriptions. For instance, “dog” is mapped to descriptions such as “a dog is barking and yipping” and “barking from a dog”. During training, categorical labels in TCC are mapped to randomly sampled descriptions from their corresponding candidate lists. These descriptions are then combined with timestamps to form TDC. Finally, we construct the audio-TCC-TDC triplets, which serve as temporally strong data.
@@ 2.2 Real Data
For real data where audio and TCC are already available, we first use LLMs to extract single-event descriptions from TCC. These descriptions are then timestamped with the TAG model [15], associating single-event descriptions with temporal occurrence to form TDC. However, we observe that the TAG model exhibits certain limitations: 1) inaccuracy when multiple events overlap; 2) occasionally fails to detect events mentioned in the TCC. Therefore, to ensure timestamp accuracy, we filter out samples with timestamp overlaps or omissions, retaining only well-curated data as temporally strong data. In addition, the original audio-TCC pairs are retained as temporally weak data and used during training to enhance generalization.
@@ 3 PicoAudio2
The overall structure of PicoAudio2 is shown in Figure 2, which comprises a Variational Autoencoder (VAE), a text encoder, and a timestamp encoder. The VAE operates on raw waveforms to compress audio signals into latents $\mathbf{A}$ during training and transforms latents back to audio waveforms during inference. We employ the VAE in EzAudio [7]. The timestamp encoder extracts temporal information from TDC, obtaining timestamp matrix $\mathbf{T}$. The Flan-T5 [17] text encoder extracts semantic information $\mathbf{C}$ from TCC. VAE and Flan-T5 are frozen during training.
Fig. 2: PicoAudio2 framework. The red arrow represents the training process while the blue represents inference. During inference, users can either provide detailed timestamps for each events, or a coarse description for LLM to infer the timestamp information.
@@ 3.1 Timestamp Encoder
The timestamp encoder extracts the timestamp matrix $\mathbf{T}$ from TDC to provide temporal information. For temporally weak data (which lack TDC), $\mathbf{T}$ is set as a fixed embedding sequence. This design enables PicoAudio2 to generate valid audio with only TCC input. For temporally strong data, the timestamp encoder first transforms single-event descriptions in TDC to event-level features $\mathbf{a}\in\mathbb{R}^{C}$ using Flan-T5. Then these features are aggregated according to timestamps of each event, resulting in $\mathbf{T}\in\mathbb{R}^{T\times C}$:
@@
For each timestep $t$, features of occurring events are summed as $\mathbf{T}_{t}$. In contrast to PicoAudio [12], which maps textual descriptions to fixed classes, PicoAudio2 is capable of handling any events described by natural language while still providing the representative timestamp matrix $\mathbf{T}$ that remains temporally aligned with the audio latent $\mathbf{A}$.
During training, TDC are obtained either through simulation or by preprocessing real data. During inference, users can provide TDC directly, or provide TCC and rely on LLMs to transform TCC to TDC. Since LLMs are sufficiently powerful to generate reasonable TDC from TCC, we focus on guiding the model to generate audio recordings that adhere to the temporal constraints specified in TDC.
@@ 4 Experiments
Table 2: General audio quality and temporal controllability results. AudioCaps-DJ (DisJoint) is the subset of AudioCaps without timestamp omissions or overlaps. Best results are in bold.
@@
PicoAudio2 39.961 2.618 12.253 0.370 3.29 0.857 0.771 4.15
Table 3: General audio quality results on the test set of AudioCaps. Best results are in bold.
[…]
AudioLDM2 [3] 22.598 2.252 9.914 0.334
Tango2 [5] 36.301 2.304 9.886 0.392
MAA2 [6] 26.699 2.335 10.392 0.372
AudioComposer [11] 27.418 2.516 7.864 0.285
PicoAudio2 27.388 2.451 12.347 0.383
Datasets
Both simulation and real data are used for training. For simulation data, we generate 64K audio clips. Each clip lasts at most 10 seconds and includes one to four events. For real data, we transform captions in AudioCaps [8] and WavCaps-ASSL subset [10] to TDC, resulting in 49K high-quality audio-TDC pairs. We also utilize around 106K temporally weak data samples from these two datasets. Overall, the total training data consists of two parts: ~106K temporally weak data (in the form of audio-TCC pairs) and ~113K temporally strong data (in the form of audio-TCC-TDC triplets). During training, a sampling ratio of 1:2 between temporally weak and strong data is used.
Experimental Setup
The time resolution of the timestamp matrix is set to 20 ms. The diffusion backbone follows EzAudio [7], with 24 layers, 16 attention heads and a hidden size of 1024. PicoAudio2 is trained for 50 epochs with a maximum learning rate of $1\times 10^{-4}$. The learning rate takes a linear decay schedule with a weight decay of $1\times 10^{-6}$. During inference, the classifier-free guidance scale is set to 7.5.
Evaluation
The evaluation is performed in terms of general audio quality and temporal controllability. For general audio quality, Frechet Distance (FD), Inception Score (IS), Kullback–Leibler divergence (KL) and CLAP score are used. For temporal controllability, we annotate the timestamps of each event using TAG [15]. As described in Section 2.2, we retain data without timestamp omissions or overlaps, naming this subset AudioCaps-DisJoint (DJ). We use Segment-F11 (Seg-F1[20] as the metric. Since the precise temporal control of multiple events is more difficult than single events, we filter AudioCaps-DJ to construct a multi-event subset (exclusively for temporal controllability evaluation), reporting the corresponding Seg-F11-ME (Multi-Event).
Besides objective metrics, we report the Mean Opinion Score (MOS): MOS-Q for audio quality and MOS-T for temporal controllability. 10 high-educated raters without hearing loss are invited to score 10 random samples from AudioCaps-DJ.
@@ 5.1 Generation Performance
We first compare PicoAudio2 with baselines, including AudioLDM2 [3], Tango2 [5], MAA2 [6] and AudioComposer [11], on general audio quality. Among them, MAA2 and AudioComposer require specified input formats to achieve temporal control so the inputs for these two models are constructed accordingly. For Tango2 and AudioLDM2, we input TCC and TDC respectively, and take the better-performing result for evaluation. Results are shown in Table 2 and 3.Per subjective and objective metrics, PicoAudio2 achieves performance comparable to leading models on both AudioCaps and AudioCaps-DJ, indicating high audio quality. On AudioCaps-DJ, AudioComposer performs notably worse. It takes both event and temporal information within its input descriptions, which may lead to suboptimal event fidelity. On AudioCaps, AudioComposer achieves better audio quality without timestamp input, further validating the negative influence of temporal information. By decoupling captions with temporal information, PicoAudio2 is capable of generating high-quality audio with either TCC or TDC inputs.
@@ 5.2 Ablation Studies
To verify the necessity of model design and data proposed in PicoAudio2, we further explore the effect of timestamp matrix and real training data. The timestamp matrix is first excluded from PicoAudio2, denoted as ‘w/o T’. This ensembles AudioComposer in that temporal information is included in the caption, so the model relies solely on cross attention with the caption feature to encode temporal information. Results in Table 2 and 3 demonstrate that the exclusion of the timestamp matrix results in inferior temporal controllability, while still maintaining reasonable audio quality, validating the effectiveness of the timestamp matrix.
@@ 6 Conclusion
Mainstream TTA models struggle to achieve high audio quality and precise temporal control on natural language descriptions. To this end, we propose PicoAudio2 in this work, incorporating new data processing pipelines and framework designs to improve temporal controllable TTA. In terms of data, PicoAudio2 designs pipelines for both simulation and real data, obtaining audio-TCC-TDC triplets for training. In terms of model, PicoAudio2 adopts a new timestamp matrix to represent event captions and their temporal occurrence to provide time-aligned feature for controllable generation. Objective and subjective evaluation show that PicoAudio2 achieves superior temporal controllability with audio quality comparable to mainstream TTA models. Due to challenges in real data annotation, PicoAudio2 is trained only on disjoint subsets of real datasets. It results in limited temporal control on overlapping events, which serves as a direction for our future work.